OpenStudy (anonymous):

Carissa has been hired as a saleswoman. She is paid a flat rate of $650 each week and earns an additional $13.25 commission for each sale. If her goal is to make $1,021 in one week, how many sales must Carissa make? a. 28 b. 50 c. 78 d. 127 @myko @aaronq @ganeshie8 @.Sam. @thomaster @uri @dan815 @RadEn

ganeshie8 (ganeshie8):

she gets $650 flat, and $13.25 for every sale she makes so total money she gets is, money = $650 + sales x $13.25

ganeshie8 (ganeshie8):

her goal is earn $1,021 1021 = 650 + sales x 13.25

ganeshie8 (ganeshie8):

solve sales

OpenStudy (anonymous):

how do i solve this equation?

ganeshie8 (ganeshie8):

to solve sales, we need to isolate it 1021 = 650 + sales x 13.25 subtract 650 both sides

ganeshie8 (ganeshie8):

1021 = 650 + sales x 13.25 -650 -650

OpenStudy (anonymous):

371

ganeshie8 (ganeshie8):

yes 371 = sales x 13.25

ganeshie8 (ganeshie8):

next, divide 13.25 both sides

ganeshie8 (ganeshie8):

371 = sales x 13.25 /13.25 /13.25 371/13.25 = sales sales = 371/13.25

OpenStudy (anonymous):

28!!!

ganeshie8 (ganeshie8):

yes ! so she has to make 28 sales to reach her goal of $1,021

OpenStudy (anonymous):

you so much. will you help me with these? Nat bought a dishwasher from a rental center for $639. He makes six monthly payments of $129.52 with his credit card. The rental center charges $2.95 for every purchase made with a credit card. He also had to pay late fees in the amount of $27 three different times. What is the total cost of the dishwasher?

OpenStudy (anonymous):

Eustis purchased a living room set for $2,784 using a 12-month deferred payment plan. The interest rate after the introductory period is 24.90%. A down payment of $225 is required as well as a minimum monthly payment of $107. What is the balance after the introductory period if only the minimum payment is made until then?

ganeshie8 (ganeshie8):

one by one :)

OpenStudy (anonymous):

your awesome thank you :)

OpenStudy (anonymous):

letsstart with nat?

ganeshie8 (ganeshie8):

yes :) Focus only on the money thats going from his pocket(credit card)

ganeshie8 (ganeshie8):

He makes six monthly payments of $129.52, and since they charge $2.95 everytime he swipes the card, so, he paid the money : 6x(129.52+2.95)

ganeshie8 (ganeshie8):

also since he got delayed in payment 3 times, he paid penalty of $27, 3 times. so overall, he paid :- 6x(129.52+2.95) + 3x27

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

ok let me try this :p

OpenStudy (anonymous):

bleh i dont know how :( i only did inside the ()

ganeshie8 (ganeshie8):

just put it in calculator

ganeshie8 (ganeshie8):

here, it is : http://www.wolframalpha.com/input/?i=6%28129.52%2B2.95%29+%2B+3%2827%29

OpenStudy (anonymous):

so which one is it? theres two answers. 875.82 and 794.82? and they are both in my answer choices :(

ganeshie8 (ganeshie8):

875.82

ganeshie8 (ganeshie8):

the other one is some other calculation. ignore it :)

OpenStudy (anonymous):

oh gotcha lol:) thank you so much! yay yay yay!

ganeshie8 (ganeshie8):

:) is the next question from financial algebra ?

OpenStudy (anonymous):

yes all of these are :( haha

ganeshie8 (ganeshie8):

omg ! no wonder these are a bit tough

ganeshie8 (ganeshie8):

Eustis purchased a living room set for $2,784 using a 12-month deferred payment plan. The interest rate after the introductory period is 24.90%. A down payment of $225 is required as well as a minimum monthly payment of $107. What is the balance after the introductory period if only the minimum payment is made until then?

OpenStudy (anonymous):

im glad im not the only one! lol. ok lets figure out eustis' situation lol

ganeshie8 (ganeshie8):

yes, living room set price is $2,784 you paid a down payment of, $225 you paid minimum monthly payments of $107 for 12 months : 12x107 so, overall, you paid how much ?

OpenStudy (anonymous):

1284?

ganeshie8 (ganeshie8):

add downpayment 225 also

OpenStudy (anonymous):

1509?

ganeshie8 (ganeshie8):

so, overall, you paid : 1284 + 225 = ?

ganeshie8 (ganeshie8):

yes u paid 1509

ganeshie8 (ganeshie8):

wats the balance amount pending ?

ganeshie8 (ganeshie8):

simply subtract 1509 from the original price of 2784

OpenStudy (anonymous):

1275?

ganeshie8 (ganeshie8):

1275 is \(\large \color{Red}{\checkmark}\)

OpenStudy (anonymous):

i wish :(

OpenStudy (anonymous):

here are the choices a. $1,990.18 b. $2,053.07 c. $3,167.18 d. $3,274.18

ganeshie8 (ganeshie8):

ok, this is compound interest problem then. u remember the formula ?

OpenStudy (anonymous):

unfortunentally not :(

ganeshie8 (ganeshie8):

ok, its like this, original price is 2784, since u made down payment of 225, the balance is, 2784-225 = 2559

ganeshie8 (ganeshie8):

after that, u made monthly payments 12 times. each month u need to calculate the interest and add it to the balance

ganeshie8 (ganeshie8):

prevously, we didnt account for the interest. BUT THE QUESTION wants you account for interest aswell. lets do it, ok ?

ganeshie8 (ganeshie8):

wats the interest ?

OpenStudy (anonymous):

.2490?

ganeshie8 (ganeshie8):

yes, thats for year. for month, the interest is .2490/12 = ?

OpenStudy (anonymous):

.02075?

ganeshie8 (ganeshie8):

so, after 1 month, balance = 2559(1+.02075) - 107

ganeshie8 (ganeshie8):

you must do it 12 times, (since you dont remember the formula)

ganeshie8 (ganeshie8):

after 1 month, balance = 2559(1+.02075) - 107 = 2505.1

ganeshie8 (ganeshie8):

after 2 months, balance = 2505.1(1+.02075) - 107 = ? ......... after 12 months, balance = ?

ganeshie8 (ganeshie8):

hope u see how complex it is.. :)

OpenStudy (anonymous):

ugh i dont know :( 2450.08?

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

not sure, im not getting one of the listed answer... @amistre64 may be able to help, when he is online..

ganeshie8 (ganeshie8):

question is this :- Eustis purchased a living room set for $2,784 using a 12-month deferred payment plan. The interest rate after the introductory period is 24.90%. A down payment of $225 is required as well as a minimum monthly payment of $107. What is the balance after the introductory period if only the minimum payment is made until then? here are the choices a. $1,990.18 b. $2,053.07 c. $3,167.18 d. $3,274.18

OpenStudy (anonymous):

right

OpenStudy (anonymous):

well thank you so much anyways you were a huge help!

ganeshie8 (ganeshie8):

np :) these are the hardest problems i feel.. . good luck !!

OpenStudy (amistre64):

Eustis purchased a living room set for $2,784 using a 12-month deferred payment plan. The interest rate after the introductory period is 24.90% A down payment of $225 is required as well as a minimum monthly payment of $107. What is the balance after the introductory period if only the minimum payment is made until then? ............................................. What defines the "introductory period", is it the 12 months defered? What is the interest rate during the introductory period? Do we assume no payments made during the introductory period? and if so, what does the minimum payment reflect? Interest only? I do not see a length of time in which to payoff the balance to zero to calculate it backwards otherwise